Circular No. 7552

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN PUPPIS T. Kato, Kyoto University, reports the discovery by Kazuyoshi Kanat-

su (Matsue, Shimane, Japan) on his T-Max 400 film exposures of a new variable star at α = 7^{h}37^{m}58^{s}, δ = −25^{o}56′51′′ (equinox 2000.0). Kanatsu finds no apparent counterpart on the Digital Sky Survey. K. Takamizawa (Saku-machi, Nagano) reports that the new object was not present on 22 T-Max 400 exposures (limiting mag 14) taken by himself between between 1994 Mar. 14 and 1999 Dec. 3; he notes the presence of a close stellar companion of mag 14.5, which is apparently a red star (red mag 13.6) in the USNO A2.0 catalogue having position end figures 58^{s}.43, 46′′.8. Available magnitudes: 2000 Jan. 15.72 UT, [11.7 (Kanatsu; T-Max 400); Nov. 28.703, 8.6 (Takamizawa; T-Max 400); Dec. 22.578, 8.8 (Takamizawa); 22.731, 8.7 (Kanatsu); 30.585, 8.8 (A. Takao, Kitakyushu, Fukuoka, Japan; unfiltered CCD); 30.612, 9.2 (R. Stubbings, Drouin, Victoria, Australia; visual).

COMET 2000 Y3 J. V. Scotti, Lunary and Planetary Laboratory, reports his discovery

of a comet with the 0.9-m Spacewatch telescope. The comet shows a coma diameter of 7′′ and a 0′.93 tail in p.a. 269^{o}; he also measures m_{2} = 19.7.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 30.16318  5^{h}56^{m}29.^{s}44  +26^{o}22′31′′.7
30.18693  5 56 28.56  +26 22 31.2  18.9
30.21016  5 56 27.63  +26 22 31.0  18.8

RZ LEONIS R. Ishioka, M. Uemura, T. Kato, H. Iwamatsu, and K. Matsumoto,

Kyoto University, report on behalf of the VSNET collaboration team: “Unfiltered and V -band CCD photometry of RZ Leo (cf. IAUC 7547) since Dec. 20.37 UT by G. Billings, G. Masi, S. Kiyota, and us shows a complex feature with amplitude ∼ 0.2 mag and period 0.07616 ± 0.00021 day (close to the quiescent photometric period of 0.0756 ± 0.0012 day; Mennickent et al. 1999, A.Ap. 352, 239) on Dec. 21 and 22 and a normal superhump feature (rapid rise; slower decline) with amplitude ∼ 0.4 mag and period 0.07857 ± 0.00022 day on Dec. 23 and 24. This suggests that the early modulations are ‘early superhumps’ observed in WZ Sge-type dwarf novae (cf. IAUC 6164, 6517) and current modulations are genuine superhumps.”

2000 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green